OpenStudy (anonymous):

Can somoene explaion how to do this? Solve: 5(x − 1) = 3(x + 2) − 7

OpenStudy (anonymous):

Use Distributive Law

OpenStudy (anonymous):

you are going to use distrubutive property

OpenStudy (highlyeducated87):

x=2 U want to explain

OpenStudy (anonymous):

@HighlyEducated87 please stop giving answers when others are tring to explain.

OpenStudy (highlyeducated87):

@lexiamee sorry :(

OpenStudy (highlyeducated87):

I can explain it out and give answer

OpenStudy (anonymous):

5(x-1) = 3(x+2) - 7 5x-5 = 3x + 6 -7 5x - 3x = 6 - 7 + 5 2x = 4 x = 2

OpenStudy (anonymous):

@ thnk you @HighlyEducated87 it is better to explain

OpenStudy (highlyeducated87):

Ok :)

OpenStudy (anonymous):

@roberto5 thank you, I get it now. @lexiamee Thanks for the start xD

OpenStudy (anonymous):

Your welcome :)